Bob's Taco Shop is considering a project that will generate a cash inflow of $11 million in year one. The project's cash flow will then grow at 5 percent per year in perpetuity.

The firm finances 25 percent of all assets with debt, and the after-tax cost of debt is 3 percent. The firm estimates the cost of equity is 15 percent.

Calculate the firm's WACC and then use the WACC to determine the maximum amount the firm should spend to undertake this project.

$6 million

$183 million

$1,100 million

$157 million

$92 million
Answer #1

WACC =(25%*3%)+(75%*15%) =12.00%

Maximum amount to be spend =Expected cash inflow/(r-g) =11/(12%-5%) =157 million